# TESSERACT: Eliminating Experimental Bias in Malware Classification across Space and Time (Extended Version)

### v1

@@ 5. Windows and PDF Malware
Comparing Android and Windows PE / PDF domains. In the realm of cybersecurity, the distinction between the Android domain and the Windows PE and PDF domains is marked and significant, especially when considering spatial and temporal aspects. The Android ecosystem, predominantly comprising mobile applications, presents a unique challenge due to its vast, diverse app market and the rapid evolution of mobile technologies. This results in a highly dynamic temporal landscape where malware characteristics and trends can shift swiftly (Zhou and Jiang, 2012; Google, 2018; Steglich et al., 2019). The diversity in the source and nature of applications, ranging from official app stores to third-party platforms, adds another layer of complexity (Felt et al., 2011). In contrast, the Windows PE and PDF domains, typically associated with desktop environments, display distinctive traits. Windows PE files, crucial for executables and libraries, are integrally linked with Windows’ architecture, leading to a steadier malware evolution, in sync with the relatively longer operation system (OS) development cycle (Ling et al., 2023; Divya and Kumar, 2016). PDF files, serving mainly for document exchange, exhibit a stable structure in format and usage (Endignoux et al., 2016; Šrndić and Laskov, 2016; Gu et al., 2022). Malicious activities via PDFs often exploit software and OS vulnerabilities. Intuitively, we could expect a slower onset of concept drift—along with a more gradual decline in model performance—in these domains.
@@
Malware ratio in the wild. The Android ecosystem predominantly features goodware (Google, 2018). Conversely, the ratio of malware to goodware in Windows PE and PDF domains is not easily quantified due to a lack of comprehensive evidence. Hence, we refer to the ratios in two popular, existing datasets: one on the EMBER dataset for Windows PE files (Anderson and Roth, 2018) (released by the EndGame company), and the other on the Hidost dataset for PDF files (Šrndić and Laskov, 2016).
5.1. Case Study: EMBER
@@
The EMBER dataset consists of a comprehensive suite of features derived from PE files to facilitate cybersecurity research (Anderson and Roth, 2018), such as headers, content-based histograms, and textual data like URL frequencies. Our analysis spans both the EMBER 2017 v2 and EMBER 2018 datasets, encompassing the years 2017 and 2018, resulting in a composite dataset of 255,499 benign and 148,231 malicious executables (labeled as having 40+ VirusTotal AV detections), constrained to this two-year period. We employ Gradient Boosted Decision Trees (GBDT) (Friedman, 2001) as our classifier, in line with the EMBER study’s methods. We initialize the classifier on data from the first three months and test it on subsequent data.
@@
The original EMBER evaluation adheres to the constraints C1 and C2 delineated in § 4.1. Defining a precise malware-to-goodware ratio for the Windows PE domain is challenging; however, the dataset’s size and heterogeneity of sources lend credence to its adherence to constraint C3. The performance degradation illustrated in 7(a) corroborates the presence of concept drift, particularly notable in the F1-score decline post-January 2018, which stabilizes at a lower benchmark. A marked reduction in the Recall metric at the onset of 2018 further underscores this point. These observations are consistent with those in the original EMBER paper (Anderson and Roth, 2018), which deliberately incorporated more elusive malware into the 2018 dataset.
@@
Applying our Tuning Algorithm on the training set improves the model’s performance on the testing set, with the overall AUT(, 21m) increasing from 0.80 to 0.83. We choose the first month data from training set to initialize the tuning model, and validate the better on the remaining two months data. We find the , compares to the original on the training data. Post-tuning, as shown in 8(a), the model displays a consistent and robust performance, maintaining a relatively stable F1-score throughout 2018. The stability was achieved by prioritizing an increase in Recall at the expense of Precision. This approach underscores the significance of establishing a realistic malware-to-goodware ratio in the training set, which is crucial for enhancing the model’s practical applicability in real-world scenarios (C3).
5.2. Case Study: Hidost
@@
Hidost, an open-source tool for extracting document structures from PDF and SWF files, forms the basis of our analysis (Šrndić and Laskov, 2016). We utilize the Hidost evaluation dataset, which includes data from a 14-week period between July 16 and October 21, 2012, comprising 1,465,737 benign and 102,234 malicious files (identified with 5+ VirusTotal AV detections). The dataset’s features are derived from static analysis of the PDF files, extracting structural paths in their hierarchy, resulting in both Boolean and numeric features. Following the original Hidost study, we employ a Random Forest (RF) algorithm (Breiman, 2001), training it on data from the initial four weeks and testing on the subsequent ten weeks.
@@
5.3. Summary of Findings
In exploring the Windows PE and PDF malware domains, we encounter the challenge of accurately defining an in-the-wild malware-to-goodware ratio, critical for training effective detection models. Our case studies on both EMBER and Hidost reveal significantly less concept drift compared to the Android dataset, likely due to the more stable and predictable nature of malware evolution in these environments, as discussed in (Chen et al., 2023b). This stability is likely linked to the slower OS development cycle in Windows PE and the consistent format and usage in PDFs. Additionally, the criteria used for labeling an instance as malware vary across different approaches and domains. For instance, in the EMBER dataset, an instance is labeled as malware if it triggers 40 or more VirusTotal alerts. Notably, we find that applying our algorithm 1 to tune the malware-to-goodware ratio during the training process enhances detection performance in testing within these domains. This also indicates the importance of further research to define this ratio accurately and to understand the mechanisms behind the benefits of tuning.
6. Delaying Time Decay
@@ 6.1. Delay Strategies
Active Learning. Active Learning (AL) strategies focus on selecting a subset of test objects, whose labels are unknown, for manual labeling and inclusion in the training set. This is based on the premise that these objects are most beneficial for updating the classification model (Settles, 2012). In our study, we employ the widely-used AL query strategy known as uncertainty sampling (Lewis and Catlett, 1994). Here, objects with the most uncertain predictions are chosen for retraining, under the assumption that they are key to refining decision boundaries. Figure 11 displays the active learning outcomes via uncertainty sampling, detailing the score (blue), Precision (yellow), and Recall (red) metrics. Notably, the data indicates discernible performance gains with the retraining of as few as 1% of samples monthly. It is observed that both Drebin and DeepDrebin achieve optimal performance when the monthly retraining rate is at its maximum of 25%, reinforcing the notion that a larger subset of retrained data is closely associated with enhanced model efficacy. It should be noted that at a monthly retraining rate of 100%, the AL approach effectively equates to Incremental Retraining.
Classification with Rejection. As a mitigation approach, classifiers can designate decisions as ”low confidence” and defer them to a later date, effectively placing uncertain predictions into a quarantine zone pending manual review (Bartlett and Wegkamp, 2008). Figure 12 demonstrates how both Drebin and DeepDrebin perform when implementing a rejection mechanism for dubious classifications, utilizing the Transcendent framework (Jordaney et al., 2017; Barbero et al., 2022). This method employs conformal evaluation theory to set thresholds that filter out the least reliable classifier decisions. Specifically, the inductive conformal evaluator (ICE) is used to pinpoint and exclude examples that deviate from expected patterns. While the original research focused on the nonconformity measure (NCM) for traditional ML algorithms, an adaptation of NCM has been developed for DNN (DeepDrebin) in this work, drawing upon the probabilistic outputs from its final SoftMax layer. Background gray histograms in the figures quantify the monthly count of quarantined objects, with cross and circle markers indicating the proportion of rejected malware and goodware, respectively.
6.2. Analysis of Delay Methods
@@
Labeling Cost (): the number of testing objects (if any) that must be labeled—the labeling must occur periodically (e.g., every month), and is particularly costly in the malware domain as manual inspection requires many resources (infrastructure, time, expertise, etc)—for example, Miller et al. (Miller et al., 2016) estimated that an average company could only manually label 80 objects per day;
@@ 7. Related Work
Experimental Bias. The base rate fallacy (Axelsson, 2000) is a well-known experimental bias in security, which highlights the limitations of TPR and FPR as performance metrics in highly imbalanced datasets. For instance, in network intrusion detection, where most traffic is benign, even a low FPR of 1% may correspond to millions of false positives and only thousands of true positives. In contrast, our work uncovers experimental settings that are misleading regardless of the metrics used, and that remain incorrect even if the correct metrics are employed (§ 4.4). While previous works such as Sommer and Paxson (Sommer and Paxson, 2010), Rossow et al. (Rossow et al., 2012), and Kouwe et al. (van der Kouwe et al., 2018) provide valuable insights and guidelines for conducting security experiments, they do not address the specific issue of temporal and spatial bias in the Android platform. Furthermore, their guidelines would not eliminate all sources of such biases as identified in our work. For example, Rossow et al. (Rossow et al., 2012) evaluate the percentage of objects in previously adopted datasets that are incorrectly labeled, without assessing the impact of such errors on classifier performance. Zhou et al. (Zhou et al., 2018) show that Hardware Performance Counters (HPCs) are not very effective for malware classification, but their focus is narrow, and they rely on 10-fold cross-validation in the evaluation. Finally, a recent paper published by Arp et al. (Arp et al., 2022) identifies ten pitfalls for machine learning in computer security, which relates to the spatio-temporal biases we discuss. The pitfalls of sampling bias and base rate fallacy overlap with spatial bias. Along with data snooping and spurious correlations relating to temporal bias. It is important to note that these connections are in part due to the paper building off our original release of Tesseract (Pendlebury et al., 2019). However, Arp et al. (Arp et al., 2022) provide the prevalence, implications, and recommendations for these pitfalls opposed to the concrete solutions provided in this work for spatio-temporal biases.
@@
Time-aware Detection. Allix et al. (Allix et al., 2015) made an important contribution by evaluating malware classifiers with respect to time and demonstrating how future knowledge can inflate performance. However, they did not propose a solution for comparable evaluations and only identified constraint C1. Additionally, Allix et al. (Allix et al., 2016a) investigated the differences between in-the-lab and in-the-wild scenarios and found that a greater presence of goodware leads to lower performance. In contrast, our work systematically analyzes and explains these issues, and addresses them by formalizing a set of constraints that jointly consider the impact of temporal and spatial bias. We introduce AUT as a unified performance metric for fair time-aware comparisons of different solutions, and offer a tuning algorithm to leverage the effects of training data distribution. While Miller et al. (Miller et al., 2016) identified temporal sample consistency (equivalent to our constraint C1), they did not identify C2 or C3, which are fundamental (§ 4.4). Moreover, they considered the test period to be a uniform time slot, whereas we take time decay into account. Roy et al. (Roy et al., 2015) questioned the use of recent or older malware as training objects and the performance degradation in testing real-world object ratios. However, their experiments were designed without considering time, which reduces the reliability of their conclusions. Past work has highlighted some sources of experimental bias (Miller et al., 2016; Roy et al., 2015; Allix et al., 2015, 2016a), but it has given little consideration to classifiers’ goals. Different scenarios may have different objectives, which may not necessarily involve maximizing . In our work, we demonstrate the effects of different training settings on performance goals and propose an algorithm to properly tune a classifier accordingly (§ 4.3). Following the publication of (Pendlebury et al., 2018, 2019), numerous studies have adopted time-aware evaluations to demonstrate their performance across various temporal stages. Notable examples include (Xu et al., 2019; Zhang et al., 2020; Barbero et al., 2022; Chen et al., 2023a).
@@
Unbiased Methods from Other Domains. The pursuit of unbiased methods is a crucial aspect in various other fields (Weiss and Provost, 2003; He and Garcia, 2009; Chawla et al., 2004). However, since these studies do not originate from the security domain, they only focus on some aspects of spatial bias and do not consider temporal bias. Concept drift is especially problematic in Android malware compared to other applications, such as image and text classification (Jordaney et al., 2017; Barbero et al., 2022). Fawcett (Fawcett, 2003) highlights challenges in spam detection, where one of the challenges is similar to spatial bias, but without providing any solution. In contrast, we propose C3 to address spatial bias and demonstrate how its violation inflates performance (§ 4.4). Torralba and Efros (Torralba and Efros, 2011) discuss dataset bias in computer vision, which is different from our security setting where there are fewer benchmarks, and the negative class (e.g., ”not cat”) in images can grow arbitrarily, which is less likely in the malware context. Moreno-Torres et al. (Moreno-Torres et al., 2012) systematize different types of drift and mention sample-selection bias, which resembles spatial bias, but they do not propose any solution or experiments to evaluate its impact on ML performance. Other related work emphasizes the importance of choosing appropriate performance metrics to avoid an incorrect interpretation of the results (e.g., ROC curves are misleading in an imbalanced dataset (Hand, 2009; Davis and Goadrich, 2006)). In this paper, we consider imbalance, propose actionable constraints and metrics, and provide tool support to evaluate the performance decay of classifiers over time.
@@
Summary. Our research is motivated by several studies on bias, but none of them comprehensively address the problem in the context of evolving data where the i.i.d. assumption does not hold. While Miller et al. (Miller et al., 2016) introduced Constraint C1, it alone is insufficient to eliminate bias. It is demonstrated by other evaluations such as (Mariconti et al., 2017), which only enforces C1. In § 4.4 of our paper, we clarify why our novel constraints C2 and C3 are fundamental and show how our AUT metric can effectively reveal the true performance of algorithms. Our approach provides counter-intuitive results that highlight the importance of considering both spatial and temporal bias when evaluating classifiers on evolving data.
8. Discussion

### v2

@@ 5 Windows and PDF Malware
Comparing Android and Windows PE / PDF domains. In the realm of cybersecurity, the distinction between the Android domain and the Windows PE and PDF domains is marked and significant, especially when considering spatial and temporal aspects. The Android ecosystem, predominantly comprising mobile applications, presents a unique challenge due to its vast, diverse app market and the rapid evolution of mobile technologies. This results in a highly dynamic temporal landscape where malware characteristics and trends can shift swiftly [77, 32, 63]. The diversity in the source and nature of applications, ranging from official app stores to third-party platforms, adds another layer of complexity [27]. In contrast, the Windows PE and PDF domains, typically associated with desktop environments, display distinctive traits. Windows PE files, crucial for executables and libraries, are integrally linked with Windows’ architecture, leading to a steadier malware evolution, in sync with the relatively longer operation system (OS) development cycle [43, 23]. PDF files, serving mainly for document exchange, exhibit a stable structure in format and usage [25, 62, 34]. Malicious activities via PDFs often exploit software and OS vulnerabilities. Intuitively, we could expect a slower onset of concept drift—along with a more gradual decline in model performance—in these domains.
@@
Malware ratio in the wild. The Android ecosystem predominantly features goodware [32]. Conversely, the ratio of malware to goodware in Windows PE and PDF domains is not easily quantified due to a lack of comprehensive evidence. Hence, we refer to the ratios in two popular, existing datasets: one on the EMBER dataset for Windows PE files [4] (released by the EndGame company), and the other on the Hidost dataset for PDF files [62].
5.1 Case Study: EMBER
@@
The EMBER dataset consists of a comprehensive suite of features derived from PE files to facilitate cybersecurity research [4], such as headers, content-based histograms, and textual data like URL frequencies. Our analysis spans both the EMBER 2017 v2 and EMBER 2018 datasets, encompassing the years 2017 and 2018, resulting in a composite dataset of 255,499 benign and 148,231 malicious executables (labeled as having 40+ VirusTotal AV detections), constrained to this two-year period. We employ Gradient Boosted Decision Trees (GBDT) [28] as our classifier, in line with the EMBER study’s methods. We initialize the classifier on data from the first three months and test it on subsequent data.
@@
The original EMBER evaluation adheres to the constraints C1 and C2 delineated in § 4.1. Defining a precise malware-to-goodware ratio for the Windows PE domain is challenging; however, the dataset’s size and heterogeneity of sources lend credence to its adherence to constraint C3. The performance degradation illustrated in 8(a) corroborates the presence of concept drift, particularly notable in the F1-score decline post-January 2018, which stabilizes at a lower benchmark. A marked reduction in the Recall metric at the onset of 2018 further underscores this point. These observations are consistent with those in the original EMBER paper [4], which deliberately incorporated more elusive malware into the 2018 dataset.
@@
Applying our Tuning Algorithm on the training set improves the model’s performance on the testing set, with the overall AUT(, 21m) increasing from 0.80 to 0.83. We choose the first month data from training set to initialize the tuning model, and validate the better on the remaining two months data. We find the , compares to the original on the training data. Post-tuning, as shown in 9(a), the model displays a consistent and robust performance, maintaining a relatively stable F1-score throughout 2018. The stability was achieved by prioritizing an increase in Recall at the expense of Precision. This approach underscores the significance of establishing a realistic malware-to-goodware ratio in the training set, which is crucial for enhancing the model’s practical applicability in real-world scenarios (C3).
5.2 Case Study: Hidost
@@
Hidost, an open-source tool for extracting document structures from PDF and SWF files, forms the basis of our analysis [62]. We utilize the Hidost evaluation dataset, which includes data from a 14-week period between July 16 and October 21, 2012, comprising 1,465,737 benign and 102,234 malicious files (identified with 5+ VirusTotal AV detections). The dataset’s features are derived from static analysis of the PDF files, extracting structural paths in their hierarchy, resulting in both Boolean and numeric features. Following the original Hidost study, we employ a Random Forest (RF) algorithm [12], training it on data from the initial four weeks and testing on the subsequent ten weeks.
@@
5.3 Summary of Findings
In exploring the Windows PE and PDF malware domains, we encounter the challenge of accurately defining an in-the-wild malware-to-goodware ratio, critical for training effective detection models. Our case studies on both EMBER and Hidost reveal significantly less concept drift compared to the Android dataset, likely due to the more stable and predictable nature of malware evolution in these environments, as discussed in [15]. This stability is likely linked to the slower OS development cycle in Windows PE and the consistent format and usage in PDFs. Additionally, the criteria used for labeling an instance as malware vary across different approaches and domains. For instance, in the EMBER dataset, an instance is labeled as malware if it triggers 40 or more VirusTotal alerts. Notably, we find that applying our algorithm 1 to tune the malware-to-goodware ratio during the training process enhances detection performance in testing within these domains. This also indicates the importance of further research to define this ratio accurately and to understand the mechanisms behind the benefits of tuning.
6 Delaying Time Decay
@@ 6.1 Delay Strategies
Active Learning. Active Learning (AL) strategies focus on selecting a subset of test objects, whose labels are unknown, for manual labeling and inclusion in the training set. This is based on the premise that these objects are most beneficial for updating the classification model [60]. In our study, we employ the widely-used AL query strategy known as uncertainty sampling [40]. Here, objects with the most uncertain predictions are chosen for retraining, under the assumption that they are key to refining decision boundaries. Figure 11 displays the active learning outcomes via uncertainty sampling, detailing the score (blue), Precision (yellow), and Recall (red) metrics. Notably, the data indicates discernible performance gains with the retraining of as few as 1% of samples monthly. It is observed that both Drebin and DeepDrebin achieve optimal performance when the monthly retraining rate is at its maximum of 25%, reinforcing the notion that a larger subset of retrained data is closely associated with enhanced model efficacy. It should be noted that at a monthly retraining rate of 100%, the AL approach effectively equates to Incremental Retraining.
Classification with Rejection. As a mitigation approach, classifiers can designate decisions as ”low confidence” and defer them to a later date, effectively placing uncertain predictions into a quarantine zone pending manual review [9]. Figure 12 demonstrates how both Drebin and DeepDrebin perform when implementing a rejection mechanism for dubious classifications, utilizing the Transcendent framework [37, 8]. This method employs conformal evaluation theory to set thresholds that filter out the least reliable classifier decisions. Specifically, the inductive conformal evaluator (ICE) is used to pinpoint and exclude examples that deviate from expected patterns. While the original research focused on the nonconformity measure (NCM) for traditional ML algorithms, an adaptation of NCM has been developed for DNN (DeepDrebin) in this work, drawing upon the probabilistic outputs from its final SoftMax layer. Background gray histograms in the figures quantify the monthly count of quarantined objects, with cross and circle markers indicating the proportion of rejected malware and goodware, respectively.
6.2 Analysis of Delay Methods
@@
Labeling Cost (): the number of testing objects (if any) that must be labeled—the labeling must occur periodically (e.g., every month), and is particularly costly in the malware domain as manual inspection requires many resources (infrastructure, time, expertise, etc)—for example, Miller et al. [48] estimated that an average company could only manually label 80 objects per day;
@@ 7 Related Work
Experimental Bias. The base rate fallacy [7] is a well-known experimental bias in security, which highlights the limitations of TPR and FPR as performance metrics in highly imbalanced datasets. For instance, in network intrusion detection, where most traffic is benign, even a low FPR of 1% may correspond to millions of false positives and only thousands of true positives. In contrast, our work uncovers experimental settings that are misleading regardless of the metrics used, and that remain incorrect even if the correct metrics are employed (§ 4.4). While previous works such as Sommer and Paxson [61], Rossow et al. [57], and Kouwe et al. [70] provide valuable insights and guidelines for conducting security experiments, they do not address the specific issue of temporal and spatial bias in the Android platform. Furthermore, their guidelines would not eliminate all sources of such biases as identified in our work. For example, Rossow et al. [57] evaluate the percentage of objects in previously adopted datasets that are incorrectly labeled, without assessing the impact of such errors on classifier performance. Zhou et al. [76] show that Hardware Performance Counters (HPCs) are not very effective for malware classification, but their focus is narrow, and they rely on 10-fold cross-validation in the evaluation. Finally, a recent paper published by Arp et al. [5] identifies ten pitfalls for machine learning in computer security, which relates to the spatio-temporal biases we discuss. The pitfalls of sampling bias and base rate fallacy overlap with spatial bias. Along with data snooping and spurious correlations relating to temporal bias. It is important to note that these connections are in part due to the paper building off our original release of Tesseract [54]. However, Arp et al. [5] provide the prevalence, implications, and recommendations for these pitfalls opposed to the concrete solutions provided in this work for spatio-temporal biases.
@@
Time-aware Detection. Allix et al. [2] made an important contribution by evaluating malware classifiers with respect to time and demonstrating how future knowledge can inflate performance. However, they did not propose a solution for comparable evaluations and only identified constraint C1. Additionally, Allix et al. [1] investigated the differences between in-the-lab and in-the-wild scenarios and found that a greater presence of goodware leads to lower performance. In contrast, our work systematically analyzes and explains these issues, and addresses them by formalizing a set of constraints that jointly consider the impact of temporal and spatial bias. We introduce AUT as a unified performance metric for fair time-aware comparisons of different solutions, and offer a tuning algorithm to leverage the effects of training data distribution. While Miller et al. [48] identified temporal sample consistency (equivalent to our constraint C1), they did not identify C2 or C3, which are fundamental (§ 4.4). Moreover, they considered the test period to be a uniform time slot, whereas we take time decay into account. Roy et al. [58] questioned the use of recent or older malware as training objects and the performance degradation in testing real-world object ratios. However, their experiments were designed without considering time, which reduces the reliability of their conclusions. Past work has highlighted some sources of experimental bias [48, 58, 2, 1], but it has given little consideration to classifiers’ goals. Different scenarios may have different objectives, which may not necessarily involve maximizing . In our work, we demonstrate the effects of different training settings on performance goals and propose an algorithm to properly tune a classifier accordingly (§ 4.3). Following the publication of [53, 54], numerous studies have adopted time-aware evaluations to demonstrate their performance across various temporal stages. Notable examples include [72, 75, 8, 14].
@@
Unbiased Methods from Other Domains. The pursuit of unbiased methods is a crucial aspect in various other fields [71, 36, 13]. However, since these studies do not originate from the security domain, they only focus on some aspects of spatial bias and do not consider temporal bias. Concept drift is especially problematic in Android malware compared to other applications, such as image and text classification [37, 8]. Fawcett [26] highlights challenges in spam detection, where one of the challenges is similar to spatial bias, but without providing any solution. In contrast, we propose C3 to address spatial bias and demonstrate how its violation inflates performance (§ 4.4). Torralba and Efros [68] discuss dataset bias in computer vision, which is different from our security setting where there are fewer benchmarks, and the negative class (e.g., ”not cat”) in images can grow arbitrarily, which is less likely in the malware context. Moreno-Torres et al. [50] systematize different types of drift and mention sample-selection bias, which resembles spatial bias, but they do not propose any solution or experiments to evaluate its impact on ML performance. Other related work emphasizes the importance of choosing appropriate performance metrics to avoid an incorrect interpretation of the results (e.g., ROC curves are misleading in an imbalanced dataset [35, 21]). In this paper, we consider imbalance, propose actionable constraints and metrics, and provide tool support to evaluate the performance decay of classifiers over time.
@@
Summary. Our research is motivated by several studies on bias, but none of them comprehensively address the problem in the context of evolving data where the i.i.d. assumption does not hold. While Miller et al. [48] introduced Constraint C1, it alone is insufficient to eliminate bias. It is demonstrated by other evaluations such as [46], which only enforces C1. In § 4.4 of our paper, we clarify why our novel constraints C2 and C3 are fundamental and show how our AUT metric can effectively reveal the true performance of algorithms. Our approach provides counter-intuitive results that highlight the importance of considering both spatial and temporal bias when evaluating classifiers on evolving data.
8 Discussion